# BadHMP: Backdoor Attack against Human Motion Prediction

### v1

@@ IV-B Global Future Sequences Poisoning
On the contrary, when poisoning the future sequences, we globally modify the 3D coordinates of all the $K$ joints. There are also two types of targets: jammed and moving.
Jammed target can be simply generated by replicating the last frame of the poisoned input sequences for the future $T$ frames. This results in the whole sequence appearing to be jammed starting from frame $N$. The poisoning process is as follows:
$\tilde{\textbf{X}}_{N+1:N+T}=\left[\tilde{\textbf{x}}_{N},\tilde{\textbf{x}}_{N},\cdots,\tilde{\textbf{x}}_{N}\right].\$ (9)
During training, the victim model efficiently learns such a distinguishing “jammed” feature from the poisoned future sequences and associates it with the backdoor trigger feature (rigid or soft) in the input sequences. Consequently, the trained victim model will output “jammed” frames whenever it encounters the backdoor trigger at the inference stage.
@@
Moving target is embedded by applying fixed trajectories to all $K$ joints. We first extract trajectories of all joints from the future sequences of the source sample, and subsequently add these trajectories to the last frame of the poisoned input sequences to generate the moving target as follows:
$\tilde{\textbf{x}}_{(j,n)}=\tilde{\textbf{x}}_{(j,N)}+\left[\textbf{x}^{\text{src}}_{(j,n)}-\textbf{x}^{\text{src}}_{(j,N)}\right],\ j\in\{1,2,\cdots,K\},$ (10)
@@
where $N+1\leq n\leq N+T$. Note that the source sample used for trajectories extraction can be either the same as or different from the one used for rigid/soft trigger generation, and scaling transformation is not required for the moving target generation.
The type of backdoor trigger and target for sample poisoning can be selected arbitrarily. Consequently, there are four possible poisoning strategies: rigid-jammed (R-J), rigid-moving (R-M), soft-jammed (S-J), and soft-moving (S-M).
Fig. 1 compares the original clean sample with the poisoned samples generated by the four different strategies, where the solid and dotted lines represent the clean and poisoned samples, respectively. The backdoor triggers are embedded only in the selected joints on the left arm, resulting in the overlapping of most parts of the clean and poisoned skeletons. For the “rigid” trigger (input sequences of Rows 1 and 2), the pose of the left arm remains unchanged throughout the input time steps. In contrast, the “soft” trigger (input sequences of Rows 3 and 4) allows the left arm to move flexibly, but along a predefined motion. Conversely, poisoned sample with the “jammed” target (output sequences of Rows 1 and 3) exhibits no movement during the output time steps, making the human skeletons appear stuck during this period. The “moving” target (output sequences of Rows 2 and 4) exhibits interframe movements, but the trajectories of all joints follow the predefined paths.
IV-C Sample Selection
Since the attacker can only poison a small fraction of the training set, it is necessary to select the top $\rho$of clean samples that are most suitable for poisoning.
Specifically, the attacker first poisons the entire clean training set to acquire the fully poisoned training set. For each pair of clean and poisoned samples, the absolute values of their differences in Acc, jerk, and BLC are calculated and weighted to measure the stealthy loss $\pazocal{L}_{s}$ of the poisoned sample as follows:
$\pazocal{L}_{s}=\lvert\text{Acc}(\textbf{X})-\text{Acc}(\tilde{\textbf{X}})\rvert$ $+\lambda_{1}\lvert\text{Jerk}(\textbf{X})-\text{Jerk}(\tilde{\textbf{X}})\rvert$ (11)
$+\lambda_{2}\lvert\text{BLC}(\textbf{X})-\text{BLC}(\tilde{\textbf{X}})\rvert,$
where $\lambda_{1}$ and $\lambda_{2}$ are positive weighting factors that balance the contributions of different terms. Finally, the top $\rho$of poisoned samples with the smallest $\pazocal{L}_{s}$ are selected to replace their counterparts in the clean training set $D_{\text{tr}}$ to generate the poisoned dataset $\tilde{D}_{\text{tr}}$.
@@ IV-C Sample Selection
TABLE I: Action-wise prediction performance of the benign and victim models on clean and poisoned test samples of the H3.6M dataset. The poisoning strategy is soft-moving (S-M), and the model architecture is LTD.
@@ V-A1 Datasets
H3.6M is the most widely used large-scale dataset for human motion prediction, comprising 3.6 million 3D human poses. It includes motion sequences of 7 actors (S1, S5, S6, S7, S8, S9, S11) performing 15 distinct actions, such as walking, eating, smoking, and phoning. The human skeleton is composed of 32 joints expressed by exponential maps. We convert these representations to 3D coordinates and use the remaining 22 joints after removing 10 redundant joints. Similar to previous works [9, 10, 18], the sequences are down-sampled to 25 frames per second (fps) with a sampling rate of 2. S5 and S11 are employed as the test and validation datasets, respectively, while the remaining 5 subjects serve as the training dataset. To balance different actions with different sequence lengths and avoid high variance, we take 256 random samples per action for test as in [10, 35].
CMU contains 8 categories of actions where the 38-joint skeletons are also originally presented by exponential maps. Like H3.6M, these presentations are converted to 3D coordinates. The training and test datasets are divided according to the configuration of [9]. CMU is evaluated on 25 joints, with 256 random samples per action.
@@ V-A2 Model Architectures
Our attack is model agnostic, which assumes that the adversary has no knowledge about the victim model’s architecture. To demonstrate the generalizability of our proposed attack, we evaluate its performance on two model architectures: LearningTrajectoryDependency (LTD) [9] and HistoryRepeatItself (HRI) [10].
@@ V-A3 Implementation Details
The model is trained to predict both short-term (0 to 500 ms) and long-term (500 to 1000 ms) future human motions. The input length $N$ is fixed at 50 frames. The output length $T$ is set to 25 for both H3.6M and CMU datasets. The default injection ratio $\rho$ is set to 10%. We use the Adam optimizer [51] and a batch size of 256 for training. The network is trained for 50 epochs, with the learning rate initially set to 0.01 and decayed by a factor of 0.96 every two epochs. For evaluation, we measure the CDE and BDE of the model at 80 and 400 ms for short-term prediction, and 560 and 1000 ms for long-term prediction.
TABLE II: Averaged CDE and BDE measured across various datasets/model architectures/poisoning strategies.
H3.6M CMU
Architecture LTD HRI Architecture LTD HRI
Model Time (ms) 80 400 560 1000 Time (ms) 80 400 560 1000 Model Time (ms) 80 400 560 1000 Time (ms) 80 400 560 1000
benign CDE 12.4 63.4 84.3 120.8 CDE 12.0 62.8 83.8 121.8 benign CDE 10.9 43.9 57.9 88.9 CDE 10.7 43.0 56.3 87.9
BDE 36.1 120.1 166.3 163.5 BDE 37.5 119.4 165.7 165.6 BDE 24.8 117.8 163.8 232.3 BDE 25.6 119.3 165.8 232.8
R-J CDE 12.4 62.9 83.3 120.8 CDE 12.2 62.9 83.7 122.1 R-J CDE 11.0 44.1 58.6 87.7 CDE 10.9 43.7 57.5 88.9
BDE 1.1 2.8 3.4 4.4 BDE 0.9 2.5 3.0 3.6 BDE 3.6 12.2 14.6 19.8 BDE 3.3 12.0 15.2 21.4
R-M CDE 12.4 62.2 81.6 117.2 CDE 12.3 62.7 83.5 120.8 R-M CDE 11.3 44.9 58.8 89.7 CDE 10.8 44.7 58.9 89.7
BDE 3.5 5.9 6.0 6.3 BDE 2.7 5.0 6.1 6.0 BDE 6.0 7.4 8.5 14.1 BDE 4.9 10.0 12.9 20.0
S-J CDE 12.4 63.3 83.9 120.6 CDE 12.1 63.3 84.7 123.3 S-J CDE 11.0 44.0 57.9 90.5 CDE 10.7 43.1 57.4 89.9
BDE 0.7 1.8 2.0 2.8 BDE 0.8 1.7 1.9 2.4 BDE 1.7 4.5 5.5 8.1 BDE 1.4 4.1 4.7 6.2
S-M CDE 12.2 62.8 83.2 119.4 CDE 11.9 61.8 82.0 119.1 S-M CDE 11.1 44.1 57.8 88.3 CDE 10.6 43.6 58.7 91.5
BDE 3.5 6.3 8.5 6.8 BDE 2.3 3.8 4.0 4.5 BDE 4.9 5.7 6.7 11.5 BDE 4.1 6.3 7.2 11.2
@@ V-B1 Fidelity and Effectiveness
To evaluate the fidelity of the poisoned model and the effectiveness of backdoor activation, we train a benign model and a victim model on $D_{\text{tr}}$ and $\tilde{D}_{\text{tr}}$, respectively, and measure their prediction performance on both clean and poisoned test sets. Table I reports the action-wise CDE and BDE of both benign and victim models, with average values in the red cells. The dataset used for the evaluation is H3.6M, and the model architecture is LTD. All poisoned training/test samples are generated by the soft-moving (S-M) poisoning strategy.
@@
Figure 2: Visualization of a victim LTD model’s prediction performance when provided with clean and poisoned input sequences. The test sample is randomly selected from the “soccer” action in the CMU dataset. Row 1: The solid line represents the clean output sequences, while the dotted line shows the victim model’s prediction on clean input sequences. Row 2: The solid line represents the poisoned output sequences, and the dotted line shows the victim model’s prediction on poisoned input sequences.
The CDE of the victim model remains very close to that of the benign model across all actions and evaluation time steps. This indicates that the “fidelity” criterion is met, as the victim model behaves like a normally trained model when processing clean test samples. However, for test samples generated by the specific poisoning strategy, the BDE of the victim model is significantly lower compared to the CDE. Additionally, the BDE accumulates much more slowly than the CDE over time. Specifically, the victim model’s average CDE and BDE are 12.2 and 3.5 at 80 ms, and 119.4 and 6.8 at 1000 ms, respectively. These results show that the embedded backdoor can be successfully activated at test time, causing the victim model to accurately produce the target sequences as intended by the attacker, even for long-term predictions. Thus, our attack also fulfills the effectiveness criterion.
The attack performances in other cases across various datasets, model architectures, and poisoning strategies are also evaluated. Due to the page limit, we only report the average CDE and BDE of these experiments in Table II. All victim models trained on the poisoned H3.6M dataset have high fidelity and the target sequence can be effectively activated, regardless of the model architectures and poisoning strategies used. On the other hand, the attack effectiveness on the CMU dataset varies across different cases. Specifically, the “soft” trigger achieves a lower BDE compared to the “rigid” trigger. We hypothesize that the feature of the “soft” trigger is more easily learned by the victim model during training. Nevertheless, the BDE measured on the CMU dataset is consistently and significantly lower than the corresponding CDE, indicating that the attack is satisfactorily effective across all cases.
@@
TABLE III: stealthiness evaluation on poisoned samples generated with (w/) and without (w/o) the sample selection step.
Dataset Metric w/o sample selection w/ sample selection
clean poisoned Diff. clean poisoned Diff.
H3.6M Acc 5.82 6.79 0.97 5.80 6.02 0.22
Jerk 6.75 7.04 0.29 5.93 5.90 0.03
BLC 236.03 240.12 4.09 239.28 240.30 1.02
CMU Acc 3.34 4.56 1.22 3.76 4.07 0.31
Jerk 5.16 7.27 2.11 5.78 6.44 0.66
BLC 235.18 242.77 7.59 235.52 239.89 4.37
@@ V-B2 Stealthiness
Table III reports the average Acc, Jerk, and BLC for paired clean and poisoned samples. The column “Diff.” shows the absolute difference between clean and poisoned samples. The 10% of clean training samples are either 1) randomly selected from the training set or 2) selected using the “sample selection” step for poisoning. The results indicate that the differences in Acc, Jerk, and BLC between clean and poisoned samples are smaller when the sample selection step is applied. Nevertheless, even without this step, randomly poisoned samples remain sufficiently stealthy. This is due to our carefully designed backdoor trigger and target, which ensure smooth and natural transitions, thereby avoiding significant changes in acceleration, jerk, and bone length.
@@ V-B3 Effect of the Injection Ratio $\rho$
The default injection ratio is set to 10% for all the experiments presented earlier. To investigate the effect of injection ratio on attack performance, we trained multiple victim LTD models on datasets poisoned with different ratios, specifically $\rho\in\{2\%,5\%,8\%,10\%,15\%\}$. The results are presented in Table IV, where $\rho=0\%$ corresponds to the benign model. The results show that the CDE of poisoned models is insensitive to changes in $\rho$. Even at an injection ratio of 15%, the CDE of the victim model remains very close to that of the benign model, thus further confirming that the victim model maintains a high fidelity even if it is heavily poisoned.
Moreover, the BDE gradually decreases as $\rho$ increases, which is expected because a higher proportion of poisoned training samples allows the victim model to better learn the association of features between the backdoor trigger and the target sequence. The 10% default injection ratio provides a good balance of effectiveness (low BDE) and stealthiness (low injection ratio) on both datasets.
TABLE IV: Attack performance under various injection ratios.
H3.6M dataset, LTD model, R-J strategy CMU dataset, HRI model, S-M strategy
@@
0 CDE 12.4 63.4 84.3 120.8 0 CDE 10.7 43.0 56.3 87.9
BDE 36.1 120.1 166.3 163.5 BDE 25.6 119.3 165.8 232.8
2 CDE 12.4 63.7 85.4 124.1 2 CDE 10.6 42.2 56.7 88.0
BDE 5.3 11.0 11.8 14.1 BDE 13.1 28.3 32.6 48.4
5 CDE 12.5 63.1 83.3 120.1 5 CDE 10.8 43.7 58.0 88.5
BDE 3.2 8.6 9.8 12.0 BDE 8.7 17.6 21.5 31.6
8 CDE 12.2 62.8 83.6 120.6 8 CDE 10.6 43.2 56.9 86.7
BDE 1.4 3.1 3.7 4.5 BDE 5.4 8.4 9.2 14.0
10 CDE 12.4 62.9 83.3 120.8 10 CDE 10.6 43.6 58.7 91.5
BDE 1.1 2.8 3.4 4.4 BDE 4.1 6.3 7.2 11.2
15 CDE 12.4 62.1 81.7 117.2 15 CDE 10.9 44.6 59.2 88.7
BDE 0.9 2.2 2.3 3.1 BDE 2.8 4.9 5.7 9.9
TABLE V: Robustness of BadHMP against fine-tuning defense.
H3.6M dataset, LTD model, R-J strategy CMU dataset, HRI model, S-M strategy
@@
Before CDE 12.4 62.9 83.3 120.8 Before CDE 10.6 43.6 58.7 91.5
BDE 1.1 2.8 3.4 4.4 BDE 4.1 6.3 7.2 11.2
After CDE 12.4 62.5 82.9 119.3 After CDE 10.8 44.0 57.6 89.1
BDE 1.6 3.9 4.5 5.7 BDE 9.0 19.5 23.3 28.8

### v2

@@ IV-B Global Future Sequences Poisoning
On the contrary, when poisoning the future sequences, we globally modify the 3D coordinates of all the $K$ joints by applying predefined trajectories. We first extract trajectories of all joints from the future sequences of the source sample, and subsequently add these trajectories to the last frame of the poisoned input sequences to generate the moving target as follows:
$\tilde{X}_{(j,n)}=\tilde{X}_{(j,N)}+\left[\overline{X}^{\text{src}}_{(j,n)}-\overline{X}^{\text{src}}_{(j,N)}\right],\ j\in\{1,2,\cdots,K\},$ (8)
@@
where $N+1\leq n\leq N+T$. Note that the source sample used for trajectories extraction can be either the same as or different from the one used for trigger embedding.
Fig. 1 visualizes the source, clean, and poisoned examples to explain the whole poisoning process. The first row represents the source sample which is randomly selected from the test set of “walking”. For the input sequences, only joints on the left arm (in green) are utilized to generated the trigger. For the output sequences, we compute the trajectories of all the $K$ joints as the target. The second and the third rows represent a clean sample of “soccer” and its poisoned version, respectively. To facilitate visual comparison, we jointly plot clean and poisoned samples in the last row. In the input time, the backdoor trigger is embedded only in the selected joints on the left arm, resulting in the overlapping of most parts of the clean and poisoned skeletons. Meanwhile, the semantic meaning of “soccer” is not damaged as we only slightly change the left arm motion of the poisoned sample. In the output time, the poisoned sequences exhibit large differences from the clean sequences because all the $K$ joints are applied with predefined trajectories extracted from the “walking” source sample.
IV-C Victim Model Poisoning
The attacker randomly selects a subset of clean samples from the training dataset and poison them. These poisoned samples can be either used to train a malicious model or released publicly as a contaminated dataset. During training, the victim model inadvertently learns a spurious correlation between the trigger (left arm motion) and the target trajectories. Consequently, when the trigger is detected during inference, the model fails to generate semantically correct predictions and instead outputting adversarial trajectories predefined by the attacker.
@@ IV-C Victim Model Poisoning
TABLE I: Action-wise prediction performance of the benign and victim LTD models on the H3.6M dataset.
@@ V-A1 Datasets
H3.6M is the most widely used large-scale dataset for human motion prediction, comprising 3.6 million 3D human poses. It includes motion sequences of 7 actors performing 15 distinct actions. The human skeleton is composed of 32 joints expressed by exponential maps. We convert these representations to 3D coordinates and use the remaining 22 joints after removing 10 redundant joints.
CMU contains 8 categories of actions where the 38-joint skeletons are also originally presented by exponential maps. Like H3.6M, these presentations are converted to 3D coordinates, and this dataset is evaluated on 25 joints.
For both two datasets, the samples are divided to training and test sets following the configuration of [9]. To balance different actions with different sequence lengths and avoid high variance, we take 256 random samples per action for testing as in [10, 38].
@@ V-A2 Model Architectures
Our attack is evaluated on two model architectures: LearningTrajectoryDependency (LTD) [9] and HistoryRepeatItself (HRI) [10].
@@ V-A3 Implementation Details
The model is trained to predict both short-term (0 to 500 ms) and long-term (500 to 1000 ms) future human motions. The input length $N$ and the output length $T$ are set to 50 and 25, respectively. The default injection ratio $\rho$ is 10%. We use the Adam optimizer and a batch size of 256 for training. The network is trained for 50 epochs, with the learning rate initially set to 0.01 and decayed by a factor of 0.96 every two epochs. For evaluation, we measure the CDE and BDE of the model at 80 and 400 ms for short-term prediction, and 560 and 1000 ms for long-term prediction.
TABLE II: Averaged CDE and BDE measured on the H3.6M dataset.
LTD HRI
Model Time (ms) 80 400 560 1000 Time (ms) 80 400 560 1000
benign CDE 12.4 62.4 81.9 118.3 CDE 11.9 62.9 84.5 123.9
BDE 37.1 121.0 166.7 168.5 BDE 36.7 119.9 168.2 170.3
victim CDE 12.1 62.2 83.0 121.1 CDE 12.2 63.0 83.7 121.1
BDE 3.1 4.8 9.2 7.1 BDE 2.7 5.3 6.7 7.4
TABLE III: Averaged CDE and BDE measured on the CMU dataset.
LTD HRI
Model Time (ms) 80 400 560 1000 Time (ms) 80 400 560 1000
benign CDE 10.8 44.4 59.6 89.8 CDE 10.6 43.0 58.0 89.1
BDE 25.4 120.2 166.1 231.5 BDE 25.6 122.3 170.3 239.5
victim CDE 10.9 44.1 58.3 88.9 CDE 10.7 43.4 57.6 87.3
BDE 6.4 5.0 5.9 8.9 BDE 5.4 4.1 5.8 7.4
@@ V-B1 Fidelity and Effectiveness
To evaluate the fidelity of the poisoned model and the effectiveness of backdoor activation, we train a benign model and a victim model on the clean training dataset $D_{\text{tr}}$ and the poisoned training dataset $\tilde{D}_{\text{tr}}$, respectively, and measure their prediction performance on both clean and poisoned test sets. Table I reports the action-wise CDE and BDE of both benign and victim models, with average values in the red cells. The dataset used for the evaluation is H3.6M, and the model architecture is LTD. All poisoned training and test samples are generated by the same source sample.
@@
The CDE of the victim model remains very close to that of the benign model across all actions and evaluation time steps. This indicates that the “fidelity” criterion is met, as the victim model behaves like a normally trained model when processing clean test samples. However, for test samples generated by the specific poisoning strategy, the BDE of the victim model is significantly lower compared to the CDE. Additionally, the BDE accumulates much more slowly than the CDE over time. Specifically, the victim model’s average CDE and BDE are 12.1 and 3.1 at 80 ms, and 121.1 and 7.1 at 1000 ms, respectively. These results show that the embedded backdoor can be successfully activated at test time, causing the victim model to accurately produce the target sequences as intended by the attacker, even for long-term predictions. Thus, our attack also fulfills the effectiveness criterion.
The attack performances in other cases across various datasets and model architectures are also evaluated. Due to the page limit, we only report the average CDE and BDE of these experiments in Table II and Table III. For all the evaluated cases, the victim model achieves low CDE comparable to that of the benign model, indicating that its prediction performance on clean test samples are not weakened. Meanwhile, the BDE measured on victim models is consistently and significantly lower than the corresponding CDE, indicating that the attack is satisfactorily effective across all cases. In summary, all victim models trained on the poisoned dataset have high fidelity and the target sequences can be effectively activated, regardless of the datasets and model architectures.
Figure 2: Visualization of a victim LTD model’s predictions on clean and poisoned input sequences. Row 1: clean output sequences (solid) and the victim model’s prediction on clean input sequences (dotted). Row 2: poisoned output sequences (solid) and the victim model’s prediction on poisoned input sequences (dotted).
@@
TABLE IV: Maximum Acc, maximum jerk, and averaged BLC measured on paired clean and poisoned training samples.
Dataset H3.6M CMU
Metric clean poisoned clean poisoned
max. Acc 44.14 43.65 13.79 13.61
max. Jerk 72.66 75.72 23.14 24.58
avg. BLC 238.87 239.08 337.19 337.35
@@ V-B2 Stealthiness
Table IV reports the maximum acceleration, maximum jerk, and averaged BLC measured on paired clean and poisoned samples.
It shows that the poisoned samples exhibit kinematically plausible motion patterns, with maximum acceleration and jerk values closely matching those of clean samples. This is a direct result of our carefully designed trigger and target, which enforce smooth spatiotemporal transitions. Meanwhile, by applying a bone-length-aware scaling transformation to the source sample before trigger and target extraction, we ensure the BLC of poisoned and clean sequences remains statistically indistinguishable. These results collectively validate the stealthiness of our attack under kinematic metrics.
@@ V-B3 Effect of the Injection Ratio $\rho$
The default injection ratio is set to 10% for all the experiments presented earlier. To investigate the effect of injection ratio on attack performance, we trained multiple victim models on datasets poisoned with different ratios, specifically $\rho\in\{2\%,5\%,8\%,10\%,15\%\}$. The results are presented in Table V, where $\rho=0\%$ corresponds to the benign model. The results show that the CDE of poisoned models is insensitive to changes in $\rho$. Even at an injection ratio of 15%, the CDE of the victim model remains very close to that of the benign model, thus further confirming that the victim model maintains a high fidelity even if it is heavily poisoned.
Moreover, the BDE gradually decreases as $\rho$ increases, which is expected because a higher proportion of poisoned training samples allows the victim model to better learn the association between the trigger and the target. The 10% default injection ratio provides a good balance of effectiveness (low BDE) and stealthiness (low injection ratio) on both datasets.
TABLE V: Attack performance under various injection ratios.
H3.6M dataset, LTD model CMU dataset, HRI model
@@
0 CDE 12.4 62.4 81.9 118.3 0 CDE 10.6 43.0 58.0 89.1
BDE 37.1 121.0 166.7 168.5 BDE 25.6 122.3 170.3 239.5
2 CDE 12.6 62.0 81.2 115.4 2 CDE 10.7 43.3 56.6 88.1
BDE 8.4 12.4 13.5 13.8 BDE 14.5 30.1 34.5 46.2
5 CDE 12.2 62.6 83.3 120.5 5 CDE 10.5 42.6 56.1 88.6
BDE 5.5 5.9 7.2 9.2 BDE 8.0 8.9 11.0 16.7
8 CDE 12.4 62.5 82.0 116.7 8 CDE 10.6 42.4 56.5 87.0
BDE 3.7 7.3 11.0 8.6 BDE 5.9 4.7 6.1 8.2
10 CDE 12.1 62.2 83.0 121.1 10 CDE 10.7 43.4 57.6 87.3
BDE 3.1 4.8 9.2 7.1 BDE 5.4 4.1 5.8 7.4
15 CDE 12.4 62.6 82.5 118.3 15 CDE 10.7 43.0 56.9 89.9
BDE 2.3 5.8 6.0 5.8 BDE 3.9 4.0 5.0 6.9
TABLE VI: Robustness of BadHMP against fine-tuning defense.
H3.6M dataset, LTD model CMU dataset, HRI model
@@
Before CDE 12.1 62.2 83.0 121.1 Before CDE 10.7 43.4 57.6 87.3
BDE 3.1 4.8 9.2 7.1 BDE 5.4 4.1 5.8 7.4
After CDE 12.0 63.1 83.9 122.7 After CDE 10.6 43.4 57.4 88.0
BDE 9.3 20.3 20.9 21.7 BDE 8.2 10.3 12.6 16.1